### v1

@@ Abstract
We consider a remote inference system with multiple modalities, where a multimodal machine learning (ML) model performs real-time inference using features collected from remote sensors. As sensor observations may change dynamically over time, fresh features are critical for inference tasks. However, timely delivering features from all modalities is often infeasible due to limited network resources. To this end, we study a two-modality scheduling problem to minimize the ML model’s inference error, which is expressed as a penalty function of AoI for both modalities. We develop an index-based threshold policy and prove its optimality. Specifically, the scheduler switches modalities when the current modality’s index function exceeds a threshold. We show that the two modalities share the same threshold, and both the index functions and the threshold can be computed efficiently. The optimality of our policy holds for (i) general AoI functions that are non-monotonic and non-additive and (ii) heterogeneous transmission times. Numerical results show that our policy reduces inference error by up to 55% compared to round-robin and uniform random policies, which are oblivious to the AoI-based inference error function. Our results shed light on how to improve remote inference accuracy by optimizing task-oriented AoI functions.
@@ I Introduction
The advent of sixth-generation (6G) technology, along with advances in artificial intelligence (AI), enables remote inference in various intelligent applications, such as autonomous transportation, unmanned mobility, and industrial automation [1]. As illustrated in Fig. 1, a remote inference system uses AI models for inference tasks (e.g., monitoring, reasoning, and decision-making) based on features transmitted from remote sensors. For instance, traffic prediction relies on near real-time forecasts of traffic status (e.g., speed, flow, and demand) based on spatio-temporal road data [2]. When the sensor observations change dynamically over time, timely data delivery is critical. For example, autonomous driving requires timely updates on vehicle positions and velocities to ensure safety, and healthcare monitoring relies on real-time vital signs for timely alerts.
Figure 1: A multimodal remote inference system.
@@
Moreover, many complex tasks involve multiple data modalities, such as audio, visual, 3D points (e.g., LiDAR or RADAR), and motion (e.g., IMU) data. Each modality provides complementary information to enhance the overall inference accuracy. Take object detection and tracking as an example: while color images (RGB) capture the shape and appearance of objects, LiDAR images offer depth information [3]. To fully exploit information from multiple modalities, machine learning (ML) based fusion techniques have been widely adopted, due to their effectiveness in extracting modality-specific information and capturing cross-modal correlations using architectures such as deep neural networks [4].
@@
Despite the advantages of multimodal ML, its application in remote inference systems remains limited in practice. One primary challenge is limited network resources (e.g., bandwidth), which makes it hard for multimodal ML models to access fresh features across all modalities simultaneously. Therefore, it is important to schedule transmissions for the modality that contributes most to the ML model’s performance. Another challenge is the lack of a clear expression capturing the relationship between feature freshness and the ML model’s performance [5]. In other words, it is unclear which feature is most valuable to the model. One main reason is the complexity of multimodal ML models, especially deep neural networks.
@@
To address these challenges, prior studies on remote inference have offered key insights into how information freshness affects ML performance [6, 7, 1]. In these studies, Age of Information (AoI), introduced in [8], serves as a crucial metric for evaluating freshness. A key finding from these works reveals that in supervised learning, ML inference error can be expressed as a function of AoI; surprisingly, this function need not be monotonic. This observation enables us to shift the goal from directly optimizing ML inference error to minimizing an AoI-based penalty function, which can be non-monotonic. Moreover, the function depends on the AoIs of all modalities and may be non-additive, i.e., not a weighted sum of single-modality AoI functions (see Section V-C for an example). The generality of this function further complicates the scheduling problem.
@@
Therefore, a key research question is the following: Under limited network resources, how can we effectively schedule modalities to minimize the ML model’s inference error?
Our main contributions are summarized as follows:
@@
  • We formulate a two-modality scheduling problem aimed at minimizing inference error, which is a general function of the AoI of both modalities. The two-modality setting (e.g., visual-audio or RGB-LiDAR) is widely used in ML, but remains underexplored in remote inference.
@@
  • We develop an optimal index-based threshold policy. Specifically, one modality is scheduled continuously until its index function exceeds a threshold, and then the scheduler switches to the other modality. Interestingly, the two modalities share the same threshold. Both the index functions and the threshold can be computed efficiently. Our results apply to (i) general AoI functions that are non-monotonic and non-additive and (ii) heterogeneous modality transmission times.
@@
  • Numerical experiments on a robot state prediction task demonstrate that our policy reduces inference error by up to 55% compared to round-robin and uniform random policies, which are agnostic to inference error. These results shed light on how to improve remote inference accuracy by optimizing task-oriented AoI functions.
@@ II Related Work
AoI penalty functions. Age of Information is a widely adopted metric for quantifying information freshness. (see [1] and references therein, as well as a comprehensive survey [9]). Nevertheless, the relationship between information freshness and its value to the application is not well understood. To address this, in [10], Sun and Cyr suggested several non-decreasing, non-linear AoI functions to capture the value of fresh data in various applications. Researchers have also proposed various metrics, such as Age of Incorrect Information (AoII) [11], to quantify the value of information in different systems; see [12] for a comprehensive survey.
@@
More recent research has explored the impact of freshness in remote inference systems. In the seminal work [6, 7], Shisher et al. demonstrated that inference error can be expressed as a general AoI penalty function, which may not be monotonic. Based on this observation, the authors studied a transmission scheduling problem to minimize a general AoI penalty function. In [13], the joint optimization of feature length selection and transmission scheduling was further studied. However, these prior works focus solely on the single-modality case, where each task corresponds to a single source. Optimizing a general AoI function for multiple modalities remains underexplored. We take a step forward by considering a general AoI function of two modalities. The new objective adds an extra layer of complexity due to the non-separability of the two modalities.
@@
Multi-source AoI optimization. There have also been efforts to optimize AoI in multi-source systems, such as queueing systems [14], broadcast networks [15], and remote inference systems [6, 7, 13]. In [14], Sun et al. studied age-optimal online scheduling in a multi-flow, multi-server queueing system, where the age function is time-dependent, symmetric, and non-decreasing. In [15], Kadota et al. demonstrated that the Maximum Age First (MAF) policy is optimal in a homogeneous network, and a suboptimal Whittle’s index method was further applied for heterogeneous scenarios. In [13], Shisher et al. proposed the Net Gain policy for jointly optimizing feature length selection and source scheduling in multi-source scenarios. Although source and modality scheduling share similarities, our work considers a more general age function. Existing works on multi-source scheduling mostly consider an additive or a non-decreasing age function, whereas we consider a potentially non-monotonic and non-additive function. Therefore, approaches such as Whittle’s index and Net Gain policy, which rely on the additive assumption, may not be directly applicable to our problem.
@@
Additionally, this paper relates to scheduling for correlated sources. Prior studies in correlated sources often model the correlation explicitly, e.g., one source may contain another’s information [16, 17]. In contrast, in our setting, the correlation among modalities is implicitly captured by the AoI-based penalty function, thus requiring a new algorithmic design.
@@ III System Model and Problem Formulation
Let $\mathbb{R}$, $\mathbb{N}$, and $\mathbb{Z}_{+}$ be the sets of real numbers, natural numbers, and positive integers, respectively.
III-A System Model
As depicted in Fig. 2, we consider a two-modality remote inference system, where two sensors send features to a receiver over a wireless channel. Each sensor corresponds to a distinct modality. Let time be slotted, indexed by $t=0,1,2,\dots$. At each time $t$, each modality $m$ generates one feature $X_{m,t}$ from a given feature set $\mathcal{X}_{m}$; the joint feature set is defined as $\mathcal{X}\coloneq\mathcal{X}_{1}\times\mathcal{X}_{2}$. Meanwhile, on the receiver side, a predictor (e.g., a neural network) uses the freshest received features from both modalities to infer the current target value $Y_{t}$ from a given target set $\mathcal{Y}$. Due to limited network resources, the scheduler can select one modality for transmission at any time. Suppose the $n$-th transmission starts at time $S_{n}$ and completes at time $D_{n}$. At time $S_{n}$, the scheduler selects only one modality for transmission, and the decision is denoted by $a_{n}\in\{1,2\}$. We assume that the scheduler always transmits the freshest feature, i.e., $X_{a_{n},S_{n}}$, from modality $a_{n}$.
Moreover, let $T_{m}\in\mathbb{Z}_{+}$ denote the fixed transmission time for modality $m$ ($T_{m}<\infty$). Because of varying feature size, $T_{1}$ and $T_{2}$ may differ. The channel is reliable, and there is no preemption during transmission. That is, the receiver successfully receives the $n$-th selected feature from modality $a_{n}$ at time $D_{n}$, and the transmission duration is $T_{a_{n}}$. When the current transmission completes, the next transmission begins immediately, i.e., $S_{n+1}=D_{n}$ for every $n$.
We use age of information to quantify information freshness, defined as the time elapsed since the freshest received feature was generated [18]. For each modality $m$, we denote its AoI at the receiver at time $t$ as $\Delta_{m}(t)\in\mathbb{Z}_{+}$. According to the definition, the AoI of modality $m$ at the receiver resets to its transmission time $T_{m}$ whenever the receiver receives a feature from modality $m$ (i.e., $a_{n}=m$ and $t=D_{n}$ for some $n$); otherwise, the AoI increases by 1. That is, the AoI of each modality $m$ evolves as
@@ III-A System Model
Figure 2: System model.
Inference error
Define
@@
We assume that the process $\{(Y_{t},X_{t}),t=0,1,2,\dots\}$ is stationary. This indicates that the inference error is time-invariant. Second, the processes $\{(Y_{t},X_{t}),t=0,1,2,\dots\}$ and $\{\mathbf{\Delta}(t),t=0,1,2,\dots\}$ are independent. This holds when the scheduling policy does not know the feature or the target value (i.e., signal-agnostic). Under these two assumptions, the expected inference error at time $t$ can be expressed as a function of the AoI vector [7]. Specifically, we use $L:\mathbb{Z}^{2}_{+}\mapsto\mathbb{R}$ to denote the expected inference error function. For every AoI vector $\boldsymbol{\delta}$, function $L(\boldsymbol{\delta})$ is defined as
@@
Assumption 1 (Uniformly Bounded $L$).
There exists a finite constant $M$ such that $|L(\boldsymbol{\delta})|\leq M$ for every AoI vector $\boldsymbol{\delta}$.
@@
Scheduling policy
The policy $\pi$ is represented as a sequence of modality choices, i.e., $\pi\coloneq(a_{0},a_{1},a_{2}\dots)$. We focus on scheduling policies that satisfy three conditions. (i) Signal-agnostic: The scheduler does not know the signal value (i.e., the feature or the target value) at any time. This is practical when the task-related signals are private to the scheduler. (ii) Causal: Each decision $a_{n}$ relies solely on the current and historical AoI at the receiver (i.e., $\mathbf{\Delta}(0),\mathbf{\Delta}(1),\dots,\mathbf{\Delta}(S_{n})$) (iii) The scheduler knows the expected inference error function $L$. In practice, this function can be numerically estimated by calculating the average loss over the dataset for each AoI vector. Let $\Pi$ denote the set of all policies satisfying these conditions.
III-B Problem Formulation
We aim to find a scheduling policy in $\Pi$ that minimizes the time-averaged expected inference error over an infinite horizon. We define this problem as Problem OPT:
@@ IV Optimal Scheduling Policy Design
The roadmap for this section is as follows. We first reformulate the problem and cast it as a Semi-Markov Decision Process (SMDP); the reformulated problem is simpler to solve. Then, we present our policy and prove its optimality. More interpretations of our policy are provided at the end.
@@ IV-A Problem Reformulation
Figure 3: An example of AoI transitions under unit transmission times ($T_{1}=T_{2}=1$). Each circle represents an AoI vector, and the number on each arrow represents the decision, i.e., the selected modality. The restart states are $(1,2)$ and $(2,1)$.
We cast the above process as a Semi-Markov Decision Process (SMDP). The process has two states $\mathbf{\Delta}_{m,\mathrm{re}},\mathbf{\Delta}_{m^{\prime},\mathrm{re}}\in\mathbb{Z}_{+}^{2}$. At state $\mathbf{\Delta}_{m,\mathrm{re}}$, the scheduler takes an action $\tau_{m}\in\mathbb{N}$. That is, it selects modality $m$ for $\tau_{m}$ consecutive times, then switches to modality $m^{\prime}$ (with one transmission from $m^{\prime}$). The transition probability from $\mathbf{\Delta}_{m,\mathrm{re}}$ to $\mathbf{\Delta}_{m^{\prime},\mathrm{re}}$ equals 1, i.e., the process switches between two states. The transition duration is given by $\tau_{m}T_{m}+T_{m^{\prime}}$. Let $C_{m}(\tau)$ denote the transition cost when the action $\tau_{m}=\tau$. For modality 1, we have
@@
which represents the total inference error during $\tau$ transmissions from modality 1 and one transmission from modality 2. Similarly, for modality 2, we have
@@
The objective is to minimize the cycle inference error, i.e.,
$\inf_{\tau_{1},\tau_{2}\in\mathbb{N}}\frac{C_{1}(\tau_{1})+C_{2}(\tau_{2})}{(\tau_{1}+1)T_{1}+(\tau_{2}+1)T_{2}}.$ (2)
Let $h(\cdot)$ denote the relative value function. The Bellman optimality equation is given by
$h(\mathbf{\Delta}_{m,\mathrm{re}})=\inf_{\tau_{m}\in\mathbb{N}}\big{[}$ $C_{m}(\tau_{m})-(\tau_{m}T_{m}+T_{m^{\prime}})\bar{L}_{\mathrm{opt}}$ (3)
@@
According to [19, Prop. 7.4.1], if the process returns to the restart state within finite time under all policies, optimizing Problem OPT is equivalent to optimizing the cycle inference error in Eq. (2). As the AoI evolution is deterministic, it suffices to find the optimal solution $\tau_{m}$. We focus on policies with finite $\tau_{m}$ for each $m$, which ensure the process returns to a restart state in finite time.
While dynamic programming methods (e.g., policy or value iteration) can solve the Bellman optimality equation (3), they suffer from the curse of dimensionality. We next present a low-complexity method to determine the optimal solution.

### v2

@@ Abstract
We consider a remote inference system with multiple modalities, where a multimodal machine learning (ML) model performs real-time inference using features collected from remote sensors. When sensor observations evolve dynamically over time, fresh features are critical for inference tasks. However, timely delivery of features from all modalities is often infeasible because of limited network resources. Towards this end, in this paper, we study a two-modality scheduling problem that seeks to minimize the ML model’s inference error, expressed as a penalty function of the Age of Information (AoI) vector of the two modalities. We develop an index-based threshold policy and prove its optimality. Specifically, the scheduler switches to the other modality once the current modality’s index function exceeds a predetermined threshold. We show that both modalities share the same threshold and that the index functions and the threshold can be computed efficiently. Our optimality results hold for general AoI functions (which could be non-monotonic and non-separable) and heterogeneous transmission times across modalities. To demonstrate the importance of considering a task-oriented AoI function, we conduct numerical experiments based on robot state prediction and compare our policy with round-robin and uniform random policies (both are oblivious to the AoI and the inference error). The results show that our policy reduces inference error by up to 55% compared with these baselines.
@@ I Introduction
The advent of sixth-generation (6G) technology, along with advances in artificial intelligence (AI), enables remote inference in various intelligent applications, such as autonomous transportation, unmanned mobility, and industrial automation [1]. As illustrated in Fig. 1, a remote inference system uses AI models for inference tasks (e.g., monitoring, reasoning, and decision-making) based on features transmitted from remote sensors. For instance, traffic prediction relies on near real-time forecasts of traffic status (e.g., speed, flow, and demand) based on spatio-temporal road data [2]. When the sensor observations change dynamically over time, timely data delivery is critical. For example, autonomous driving requires timely updates on vehicle positions and velocities to ensure safety; healthcare monitoring relies on real-time vital signs for timely alerts.
Figure 1: A multimodal remote inference system
@@
Moreover, such complex tasks often involve multiple data modalities, such as audio, visual, 3D points (e.g., LiDAR or RADAR), and motion (e.g., IMU) data. Each modality provides complementary information to enhance the overall inference accuracy. Take object detection and tracking as an example: while color images (RGB) capture the shape and appearance of objects, LiDAR images offer depth information [3]. To fully exploit information from multiple modalities, machine learning (ML) based fusion techniques have been widely adopted, as they can effectively extract modality-specific information and capture cross-modal correlations using architectures such as deep neural networks [4].
@@
Despite the advantages of multimodal ML, its application in remote inference systems remains limited in practice. One primary challenge is limited network resources (e.g., bandwidth), which makes it hard for multimodal ML models to access fresh features across all modalities simultaneously. Therefore, it is important to schedule transmissions for the modality that contributes most to the ML model’s performance. Another challenge is the lack of an accurate model that captures the relationship between feature freshness and inference accuracy [5]. That is, it is unclear which feature, at which time, is most valuable to the model. One main reason is the complexity of multimodal ML models, especially deep neural networks.
@@
To address these challenges, prior studies on remote inference have offered key insights into how information freshness affects ML performance [6, 7, 1]. In these studies, Age of Information (AoI), introduced in [8], serves as a crucial metric for measuring freshness. A key finding from these works reveals that in supervised learning, ML inference error can be expressed as a (possibly non-monotonic) function of AoI. This observation enables us to shift the goal from directly optimizing ML inference error to minimizing an AoI-based penalty function (which could be non-monotonic). Moreover, the function depends on the AoI of each modality and may be non-additive (i.e., not a weighted sum of single-modality AoI functions); see Section V-B for an example. The generality of this function further complicates the scheduling problem.
@@
To that end, in this paper, we study a key research question: How to effectively schedule transmissions for multiple modalities to minimize the ML model’s inference error under the network bandwidth constraint? We answer this question by making the following main contributions:
@@
  • We formulate a two-modality scheduling problem that seeks to minimize inference error, which is a general function of the AoI of both modalities. The two-modality setting (e.g., visual-audio or RGB-LiDAR) is common in ML, but remains underexplored in remote inference.
@@
  • We develop an optimal index-based threshold policy. Specifically, one modality is scheduled continuously until its index function exceeds a predetermined threshold, and then the scheduler switches to the other modality. Interestingly, the two modalities share the same threshold. Both the index functions and the threshold can be computed efficiently. Our results hold for general AoI functions (which could be non-monotonic and non-separable) and heterogeneous transmission times across modalities.
@@
  • We conduct numerical experiments based on robot state prediction to demonstrate the importance of optimizing a general function of AoI. The results show that our policy reduces inference error by up to 55% compared with round-robin and uniform random policies (both are agnostic to the AoI and the inference error). We believe that these results provide valuable insights into improving the accuracy of multimodal remote inference through the optimization of task-oriented AoI functions.
@@ II Related Work
AoI penalty functions. Age of Information (AoI) is a widely adopted metric for quantifying information freshness (see [1] and references therein, as well as a comprehensive survey [9]). Nevertheless, the relationship between information freshness and its value to the application is still not well understood. To that end, in [10], Sun and Cyr suggested several non-decreasing, non-linear AoI functions to capture the value of fresh data in various applications. Researchers have also proposed various metrics, such as Age of Incorrect Information (AoII) [11], to quantify the value of information in different systems (see [12] for a comprehensive survey).
@@
More recent research has explored the impact of freshness in remote inference systems. In the seminal work [6, 7], Shisher et al. demonstrated that inference error can be expressed as a general AoI penalty function, which may not be monotonic. Based on this observation, the authors studied a transmission scheduling problem, aiming to minimize a general AoI penalty function. In [13], the joint optimization of feature length selection and transmission scheduling was further studied. These prior works assume that each task corresponds to a single source, which can be viewed as a single-modality setting. Optimizing a general AoI function for multiple modalities remains underexplored. We take a step further by considering a general AoI function of two modalities, which is non-separable and thus adds an extra layer of complexity to the problem.
@@
Multi-source AoI optimization. There have also been efforts to optimize AoI in multi-source systems, such as queueing systems [14], broadcast networks [15], and remote inference systems [6, 7, 13]. In [14], Sun et al. studied age-optimal online scheduling in a multi-flow, multi-server queueing system, where the age function is time-dependent, symmetric, and non-decreasing. In [15], Kadota et al. showed that the Maximum Age First (MAF) policy is optimal in a homogeneous network, and a suboptimal Whittle’s index method was further applied for heterogeneous scenarios. In [13], Shisher et al. proposed the Net Gain policy for jointly optimizing feature length selection and source scheduling in multi-source scenarios. Although source and modality scheduling share similarities, our work considers a more general age function. Existing works on multi-source scheduling consider an additive or a non-decreasing age function, whereas a non-monotonic and non-additive age function naturally arises in the multimodal setting we consider. Therefore, approaches such as Whittle’s index and Net Gain policy, which rely on the additive assumption, may not be directly applicable to our problem.
@@
Additionally, scheduling for correlated sources is related to our work. Prior studies in correlated sources often model the correlation explicitly (e.g., one source may contain another’s information) [16, 17]. In contrast, in our setting, the correlation among modalities is implicitly captured by the AoI-based penalty function, thus requiring a new algorithmic design.
@@ III System Model and Problem Formulation
In this section, we model the multimodal remote inference system and formulate a two-modality scheduling problem.
As depicted in Fig. 2, we consider a multimodal remote inference system, where two sensors send features to a receiver over a wireless channel. Each sensor corresponds to a distinct modality $m\in\{1,2\}$. Let time be slotted, indexed by $t=0,1,\dots$. At each time $t$, each modality $m$ generates one feature $X_{m,t}$ from a given feature set $\mathcal{X}_{m}$; the joint feature set is defined as $\mathcal{X}\coloneq\mathcal{X}_{1}\times\mathcal{X}_{2}$. Meanwhile, on the receiver side, a predictor (e.g., a neural network) uses the freshest received features from both modalities to infer the current target value $Y_{t}$ from a given target set $\mathcal{Y}$. Because of bandwidth constraints, the scheduler can only select one modality for transmission at any given time. Suppose the $n$-th transmission starts at time $S_{n}$ and completes at time $D_{n}$. At time $S_{n}$, the scheduler selects only one modality for transmission, and the decision is denoted by $a_{n}\in\{1,2\}$. We assume that the scheduler always transmits the freshest feature, i.e., $X_{a_{n},S_{n}}$, from modality $a_{n}$.
Moreover, let $T_{m}$ denote the fixed transmission time for modality $m$, which is a positive finite integer. Since the feature size varies, $T_{1}$ and $T_{2}$ may differ. We assume a reliable channel and do not consider preemption during transmission. That is, the receiver successfully receives the $n$-th selected feature from modality $a_{n}$ at time $D_{n}$, and the transmission duration is $T_{a_{n}}$. We also assume a work-conserving system: when the current transmission completes, the next transmission begins immediately, i.e., $S_{n+1}=D_{n}$ for every $n$.
We use age of information (AoI) to quantify information freshness, defined as the time elapsed since the freshest received feature was generated [18]. For each modality $m$, we denote its AoI at the receiver at time $t$ as $\Delta_{m}(t)\in\mathbb{Z}_{+}$, where $\mathbb{Z}_{+}$ denotes the set of positive integers. According to the definition, the AoI of modality $m$ at the receiver resets to its transmission time $T_{m}$ whenever the receiver receives a feature from modality $m$ (i.e., $a_{n}=m$ and $t=D_{n}$ for some $n$); otherwise, the AoI increases by 1. That is, the AoI of each modality $m$ evolves as
@@
Figure 2: System model
Inference error. Define
@@
We assume that the process $\{(Y_{t},X_{t}),t=0,1,\dots\}$ is stationary. This implies that the inference error is time-invariant. Second, the processes $\{(Y_{t},X_{t}),t=0,1,\dots\}$ and $\{\mathbf{\Delta}(t),t=0,1,\dots\}$ are independent. This holds when the scheduling policy does not know the feature or the target value (i.e., signal-agnostic). Under these two assumptions, the expected inference error at time $t$ can be expressed as a function of the AoI vector [7]. We use $L:\mathbb{Z}^{2}_{+}\mapsto\mathbb{R}$ to denote the expected inference error function, where $\mathbb{R}$ is the set of real numbers. For every AoI vector $\boldsymbol{\delta}$, function $L(\boldsymbol{\delta})$ is defined as
@@
Assumption 1.
There exists a finite constant $M$ such that $|L(\boldsymbol{\delta})|\leq M$ for all AoI vectors $\boldsymbol{\delta}\in\mathbb{Z}^{2}_{+}$.
@@
Scheduling policy. The policy $\pi$ is represented as a sequence of modality choices, i.e., $\pi\coloneq(a_{0},a_{1},a_{2},\dots)$. We focus on scheduling policies that satisfy three conditions. (i) Signal-agnostic: The scheduler does not know the signal value (i.e., the feature or the target value) at any time. This is practical when these signals are private to the scheduler. (ii) Causal: Each decision $a_{n}$ relies solely on the current and historical AoI (i.e., $\mathbf{\Delta}(0),\mathbf{\Delta}(1),\dots,\mathbf{\Delta}(S_{n})$) (iii) The scheduler knows the expected inference error function. In practice, this function can be numerically estimated by calculating the average loss over the dataset for each AoI vector. Let $\Pi$ denote the set of all policies satisfying these conditions.
Objective. We aim to design a scheduling policy in $\Pi$ that minimizes the time-averaged expected inference error over an infinite horizon. We define this problem as Problem OPT:
@@ IV Optimal Scheduling Policy Design
The roadmap for this section is as follows. We first reformulate the problem and cast it as a Semi-Markov Decision Process (SMDP), as the reformulated problem is simpler to solve. Then, we present our policy, prove its optimality, and offer interpretations of our policy.
@@ IV-A Problem Reformulation
Figure 3: AoI transition during the $i$-th cycle under unit transmission times ($T_{1}=T_{2}=1$). The process has two states: $(1,2)$ and $(2,1)$. Each pair represents an AoI vector, and the number on each arrow represents the selected modality.
In the reformulated SMDP, the policy is a sequence of decisions made at each restart state. Specifically, define a cycle as the period from state $\mathbf{\Delta}_{1,\mathrm{re}}$ back to itself. The policy is defined as $\pi_{\tau}\coloneq((\tau_{1,0},\tau_{2,0}),(\tau_{1,1},\tau_{2,1}),\dots)$, where $\tau_{m,i}$ is the decision for state $\mathbf{\Delta}_{m,\mathrm{re}}$ in the $i$-th cycle. Fig. 3 illustrates one cycle of the process in a special case of unit transmission times (i.e., $T_{1}=T_{2}=1$). As shown in Fig. 3, the new policy $\pi_{\tau}$ is derived from the original policy $\pi$ by grouping transmissions for the same modalities. Although the original policy observes more AoI states than the new policy (such as $(1,3),\dots,(1,2+\tau_{1,i})$), it does not help the scheduling, as the AoI evolution is deterministic given decisions. Hence, the two policies are equivalent, and Problem OPT is equivalent to
$\inf_{\pi_{\tau}\in\Pi}\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{\pi_{\tau}}\left[\sum_{t=0}^{T-1}L\left(\mathbf{\Delta}(t)\right)\right].$ (2)
According to [19, Proposition 11.4.7.], if the state space is finite and the process always returns to each state within a finite time, then solving Problem (2) is equivalent to solving the Bellman optimality equation. To ensure this, we consider policies that switch to each modality in finite time; that is, there exists a finite number $\tau_{\max}\in\mathbb{N}$ such that $\tau_{m,i}\leq\tau_{\max}$ for all $m$ and $i$.
We next present the Bellman optimality equation. For each $m$, the transition probability from $\mathbf{\Delta}_{m,\mathrm{re}}$ to $\mathbf{\Delta}_{m^{\prime},\mathrm{re}}$ equals 1. The transition duration is $\tau T_{m}+T_{m^{\prime}}$. Let $C_{m}(\tau)$ denote the transition cost when the decision is $\tau$. For modality 1, we have
@@
which represents the total inference error during $\tau$ transmissions for modality 1 and one transmission for modality 2. Similarly, for modality 2, we have
@@
The Bellman optimality equation is given by
$h(\mathbf{\Delta}_{m,\mathrm{re}})=\min_{\tau\in\{0,1,\dots,\tau_{\max}\}}\big{[}$ $C_{m}(\tau)-(\tau T_{m}+T_{m^{\prime}})\bar{L}_{\mathrm{opt}}$ (3)
@@
While dynamic programming methods (e.g., policy or value iteration) can solve the Bellman optimality equation (3), they suffer from the curse of dimensionality. Therefore, we aim to design a low-complexity policy that is optimal.